complete the square of y = ax 2+bx+c
Thank You:)

Answer by stanbon(75887)

complete the square of y = ax^2+bx+c
--------------------
ax^2 + bx + c = 0
----
x^2 + (b/a)x + (c/a) = 0
----
x^2 + (b/a)x + (b/2a)^2 = -(c/a) + (b/2a)^2
-----
x^2 + (b/a)x + (b/2a)^2 = -(c/a) + b^2/4a^2
----
x^2 + (b/a)x + (b/2a)^2 = (-4ac + b^2)/(4a^2)
-----
Take the square root of both sides to get:
x + (b/2a) = +/-sqrt(b^2 - 4ac)/(2a)
----
x = [-b +- sqrt(b^2-4ac)]/(2a)
========================================
Cheers,
Stan H.
============